John wants to buy a new sports car, and he estimates that he'll need to make a $3,025.00 down payment towards his purchase. If he has 17 months to save up for the new car, how much should he deposit into his account if the account earns 3.979% compounded continuously so that he may reach his goal? John needs to deposit how much money?
Answer by Theo(13342)

the formula for continuous compounding is f = p * e^(rt)

f is the future value
p is the present value
r is the interest rate per time period
t is the number of time periods
e is the scientific constant 2.718281828.....

in your problem:

p = what you want to find
f = 3025
r = 3.979% / 100 = (.03979/12) = per month
t = 17 months

formula becomes 3025 = p * e^(.03979/12 * 17)

solve for p to get p = 3025 / (e^(.03979/12 * 17) = 2859.20017.